AAVSOnet status, 2016-05-23

Mon, 05/23/2016 - 13:31

BSM_S operating nominally.  I'm holding off processing recent data until we get a -20C bias frame from the ST10 that was used during Feb/Mar.  Once those last half-dozen nights are processed, then more recent BSM_S data will be processed.

BSM_Berry operating nominally.

BSM_NM  operating nominally, now that I've figured out how to use the EQ-6 mount.

BSM_HQ operating nominally.

BSM_Hamren off-line until we replace the control computer.  Doc Kinne, Helmar Adler and myself worked on the  Microsoft Surface Pro tablet, and it looks like this will be a reasonable low-power alternative.  An anonymous donor has contributed mucn-needed funds to pay for the tablet.  With a little more work, we should be sending the tablet to Mike Linnolt later this week for installation.

coker30 operating nominally.

SRO operating nominally.

OC61 operating nominally.

TMO61  operational, but need to get approval from Jon Holtzman to use the system.  We hope this will be back on-line in a few days.

The southwestern monsoon will be starting over the next month or so.  It usually hits New Mexico first and then moves westward, so TMO61 and BSM_NM are likely to be impacted soon and will then be off-line until September.